OpenStudy (anonymous):

At sea level 3978(mi from center of earth) an astronaut weighs 117lbs. Find the weight when it is 284 mi above surface of the earth & the spacecraft is not in motion.

OpenStudy (anonymous):

can you provide a relevant equation maybe?

OpenStudy (anonymous):

formula*

OpenStudy (anonymous):

W=k/d^2

OpenStudy (anonymous):

117=k/3978^2

OpenStudy (anonymous):

k=1851464628

OpenStudy (anonymous):

what is k?

OpenStudy (anonymous):

w = work i presume?

OpenStudy (anonymous):

shouldn't you use the universal gravitational formula for this problem? (haven't taken physics for a little)

OpenStudy (anonymous):

w=weight

OpenStudy (anonymous):

no, weight is just mass times gravity w=mg

OpenStudy (anonymous):

I am asking for help from my professor this is what she is giving me

OpenStudy (anonymous):

what's k in the formula

OpenStudy (anonymous):

not the number, what does it stand for?

OpenStudy (anonymous):

k=just a variable i think

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

k = 1851464628

OpenStudy (anonymous):

w = 1851464628/284^2

OpenStudy (anonymous):

booya there ya'go i think

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

4262^2*

OpenStudy (anonymous):

I got 101.9

OpenStudy (anonymous):

which was correct